# Scout-Rover cooperation: online terrain strength mapping and traversal risk estimation for planetary-analog explorations

### v2

@@ 1 Introduction
To address this issue, we propose to use high-mobility legged robots as scouts to directly assess regolith properties through locomotion and complement rover operations. Legged robots are typically limited in payload capacity as compared to wheeled rovers, but could offer advantages in traversing deformable and rugged terrains  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?). In addition, their physical interaction with the terrain during each step provides a unique and versatile sensing modality: every footfall generates direct measurements of ground properties  (?, ?, ?, ?, ?). Recent advances in actuator technology — particularly in direct-drive and quasi-direct-drive actuators with minimal gearing backlash — has enabled high actuator transparency and proprioceptive precision  (?, ?). This allows legged robots to accurately measure ground reaction forces and infer terrain mechanics  (?, ?, ?). Unlike visual-based terrain assessment  (?, ?, ?, ?, ?), which can struggle to detect deformable or subsurface hazards, interaction-based sensing provides rich information about the mechanical properties of homogeneous and heterogeneous granular terrain, including penetration resistance, shear strength, layering, and cohesion  (?, ?, ?). These directly-sensed terrain parameters are critical for evaluating the traversal risk of different rover configurations and for informing safe and efficient path planning. By pairing legged scouts with wheeled rovers, we leverage the complementary strengths of both platforms: legged robots provide dense, physics-grounded terrain measurements, while wheeled rovers execute payload-intensive scientific operations along routes informed by those measurements. This cooperative paradigm enables terrain-aware operations that expand the reachable science workspace without sacrificing the robustness and efficiency of wheeled platforms.that improve the rover mission safety and efficiency.
@@ 2.2.1 Regolith Preparation and Strength Ground Truthing
We deployed our regolith mapping and traversal risk estimation module in a 6-meter by 4-meter section of the Ames SSERVI Regolith Lab. The testbed was filled with LHS-1 lunar simulant to an average depth of 7–8 cm (Fig. 2A) to minimize bottom boundary effects that can influence force transmission in shallow granular beds  (?). To generate controllable variation in regolith strength and resulting variation in traversal risk, the simulant was divided into a 1-meter grid (Fig. 2E, F). Each cell was manually prepared to one of three compaction levels by applying distinct surface treatments: (1) tamping, which prepares regolith to high compaction and strength by firmly compacting the simulant using a hand tamper (Fig. 2B); (2) raking, which prepares regolith to medium compaction and strength by loosening and evenly redistributing the simulant surface using a garden rake (Fig. 2C); and (3) sifting, which prepares the simulant to low compaction and strength by passing them through a mesh sieve (Fig. 2D). Using these three treatments, the testbed was prepared to a pre-designed spatial variation pattern of regolith strength as shown in Fig. 2E, F. After the regolith was prepared following the predesigned pattern, the surface appeared relatively flat, with height standard deviations of 1.66 centimeters according to the LIDAR scan (Fig. 2I). Visually, it was difficult to discern the differences in terrain traversability. To determine the strength across the testbed, we collected ground-truth regolith strength measurements using a robotic leg penetrometer, Traveler (Fig. 2G). The Traveler toe penetrates vertically into the regolith, while measuring how quickly the ground reaction forces increases with the penetration depth (see Sec. 4.3). The slope of ground reaction force with respect to depth – referred to as penetration resistance – is a commonly-used metric for assessing regolith normal strength  (?). The accuracy of the robotic leg penetrometer has been extensively validated in previous studies, and deployed in a wide range of laboratory and field missions to characterize regolith properties  (?, ?, ?, ?). Leg penetrometer measurements shown that the three treatments yielded well-separated resistance values, where the penetration resistance per unit area of the tamped, raked, and sifted terrains was measured as 5.07 $\pm$ 1.79 N/cm3, 1.34 $\pm$ 0.35 N/cm3, and 0.51 $\pm$ 0.47 N/cm3, respectively (Fig. 2K). Despite small variability due to manual terrain preparation, the measured resistance values clearly distinguished the three terrain conditions and confirmed the intended distribution pattern (Fig. 2J).
@@ 2.2.3 Traversal Risk Estimation Results and Simulation Validation
To validate our risk prediction framework, we conducted comprehensive simulation tests using the Project Chrono open-source physics engine  (?), which has been systematically validated against real-world experiments  (?). We created a virtual replica of the NASA Ames LHS-1 testbed by using the ground-truth regolith strength measurements collected by the Traveler leg penetrometer to parameterize the simulation environment. Specifically, we computed the average penetration resistance value for each 1 m $\times$ 1 m cell and mapped these values to corresponding elastic and damping parameters for the SCM (Soil Contact Model) terrain representation used in the simulation  (?) (see Appendix Appendix A: Simulation Parameter Mapping for the complete stiffness parameter matrix and mapping methodology). This spatial parameterization creates a 18m $\times$ 12m grid where each cell corresponds to a 3 m $\times$ 3 m meter region of the physical testbed, with penetration resistance per unit area ranging from 0.28 to 8.38 N/cm³ and reflecting the three distinct terrain types (tamped, raked, and sifted), represented by different colors in the simulation visualization (Fig. 4, bottom section). Using this simulation environment, we tested three critical scenarios that demonstrate the importance of payload-aware risk assessment.
@@ 2.3.1 Site selection and field workflow
To demonstrate the advantages of using a scout robot to support scientific target selection and safe rover traversal, we deployed our scout–rover cooperation system at White Sands National Park. White Sands is an active gypsum dune field within a playa–lake basin that displays a rich variety of sedimentary features and surface textures, including loose sand, partially lithified sandstones, and salt and microbial crusts. Features at this site closely resemble the terrains encountered at Meridiani Planum by the NASA Opportunity rover (?, ?, ?), and the existing robust body of work characterizing this dune field facilitates our investigations (cite).
@@ 2.3.3 Terrain-aware rover navigation based on scouting measurements
The integration of risk maps and selected scientific interest goals provides rich information for mission planning that allows rovers to safely navigate to scientific targets while avoiding high-risk areas. To demonstrate this capability, we used a reactive potential field planner  (?, ?), in which goal locations corresponding to high scientific reward generate attractive forces that pull rovers closer and the boundaries of high-risk regions generate repulsive forces that push rovers away from hazardous terrain. The path that the rover takes is then dictated by the sum of attractive and repulsive forces acting at the rover at that each point in time. This approach ensures that the rover navigates along a path that achieves scientific discovery goals while minimizing the probability of immobilization. Note that for the purposes of path planning, the rover was modeled as a point mass with a bounded acceleration.
@@ 4.4 Regolith strength map reconstruction
To predict and map the terrain properties over a continuous 2D area, we use Gaussian Process Regression  (?) (GPR), also known as kriging, thereby modeling the underlying terrain property distribution (Fig. 7G, H). By updating the the predicted distributions for terrain property map and fit error heuristic map with each step, we provide real-time prediction and quantifiable uncertainty for the terrain properties and the “trustworthiness” across the landscape for planning traversal and science objectives. To use GPR, we define the kernel as
@@ 4.7 Autonomous mission execution through potential-field-based navigation
Figure 2: Experimental setup for terrain-aware legged robot scouting. (A) Experiment site, the SSERVI Regolith Lab testbed located at the NASA Ames Research Center. (B–D) Three terrains preparation protocols with strength from high to low: tamped (high strength), raked (medium strength), and sifted (low strength). (E)(F) The desired terrain strength map, prepared using the three protocols. (G) Ground truth terrain strength measurements are collected using a robotic penetrometer across the testbed. (H) A legged scout robot performs terrain scouting using proprioception from every step along its path. (I) LiDAR scan after terrain preparation. (J) Penetration resistance measurements from the robotic penetrometer across the testbed. (K) The penetration resistance distribution from terrain grid cells prepared using the three protocols.
@@
Figure 5: Workflow for safe rover mission planning integrating satellite imagery, in-situ measurements, and science-interest-driven goal selection. (A) Overview of the mission plan showing a top-down view of the operational area and the tentative traversal region (purple rectangle). (B) LiDAR-based reconstruction of the traversal region with elevation data (1983–1988 m). (C) LiDAR-based inclination measurements, and initial identification of candidate rover routes. (D) Field interface enabling real-time visualization of terrain strength and science goal selection: (a) estimated toe force over time; (b) toe force during the penetration phase; (c) model fitness; (d) penetration resistance measured from each step; and (e) interpolated penetration resistance map generated via Gaussian Process regression. (E) Field snapshot after Spirit robot scouting. (F) Final penetration resistance map of the scouted region.

### v3

@@ 1 Introduction
To address this issue, we propose to use high-mobility legged robots as scouts to directly assess regolith properties through locomotion and complement rover operations. Legged robots are typically limited in payload capacity as compared to wheeled rovers, but could offer advantages in traversing deformable and rugged terrains  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?). In addition, their physical interaction with the terrain during each step provides a unique and versatile sensing modality: every footfall generates direct measurements of ground properties  (?, ?, ?, ?, ?). Recent advances in actuator technology — particularly in direct-drive and quasi-direct-drive actuators with minimal gearing backlash — has enabled high actuator transparency and proprioceptive precision  (?, ?). This allows legged robots to accurately measure ground reaction forces and infer terrain mechanics  (?, ?, ?). Unlike visual-based terrain assessment  (?, ?, ?, ?, ?), which can struggle to detect deformable or subsurface hazards, interaction-based sensing provides rich information about the mechanical properties of homogeneous and heterogeneous granular terrain, including penetration resistance, shear strength, layering, and cohesion  (?, ?, ?). These directly-sensed terrain parameters are critical for evaluating the traversal risk of different rover configurations and for informing safe and efficient path planning. By pairing legged scouts with wheeled rovers, we leverage the complementary strengths of both platforms: legged robots provide dense, physics-grounded terrain measurements, while wheeled rovers execute payload-intensive scientific operations along routes informed by those measurements. This cooperative paradigm enables terrain-aware operations that expand the reachable science workspace without sacrificing the robustness and efficiency of wheeled platforms.
@@ 2.2.1 Regolith Preparation and Strength Ground Truthing
We deployed our regolith mapping and traversal risk estimation module in a 6-meter by 4-meter section of the Ames SSERVI Regolith Lab. The testbed was filled with LHS-1 lunar simulant to an average depth of 7–8 cm (Fig. 2A) to minimize bottom boundary effects that can influence force transmission in shallow granular beds  (?). To generate controllable variation in regolith strength and resulting variation in traversal risk, the simulant was divided into a 1-meter grid (Fig. 2E, F). Each cell was manually prepared to one of three compaction levels by applying distinct surface treatments: (1) tamping, which prepares regolith to high compaction and strength by firmly compacting the simulant using a hand tamper (Fig. 2B); (2) raking, which prepares regolith to medium compaction and strength by loosening and evenly redistributing the simulant surface using a garden rake (Fig. 2C); and (3) sifting, which prepares the simulant to low compaction and strength by passing them through a mesh sieve (Fig. 2D). Using these three treatments, the testbed was prepared to a pre-designed spatial variation pattern of regolith strength as shown in Fig. 2E, F. After the regolith was prepared following the predesigned pattern, the surface appeared relatively flat, with height standard deviations of 1.66 centimeters according to the LIDAR scan (Fig. 2I). Visually, it was difficult to discern the differences in terrain traversability. To determine the strength across the testbed, we collected ground-truth regolith strength measurements using a robotic leg penetrometer, Traveler (Fig. 2G). The Traveler toe penetrates vertically into the regolith, while measuring how quickly the ground reaction forces increase with the penetration depth (see Sec. 4.3). The slope of ground reaction force with respect to depth – referred to as penetration resistance – is a commonly-used metric for assessing regolith normal strength  (?). The accuracy of the robotic leg penetrometer has been extensively validated in previous studies, and deployed in a wide range of laboratory and field missions to characterize regolith properties  (?, ?, ?, ?). Leg penetrometer measurements shown that the three treatments yielded well-separated resistance values, where the penetration resistance per unit area of the tamped, raked, and sifted terrains was measured as 5.07 $\pm$ 1.79 N/cm3, 1.34 $\pm$ 0.35 N/cm3, and 0.51 $\pm$ 0.47 N/cm3, respectively (Fig. 2K). Despite small variability due to manual terrain preparation, the measured resistance values clearly distinguished the three terrain conditions and confirmed the intended distribution pattern (Fig. 2J).
@@ 2.2.3 Traversal Risk Estimation Results and Simulation Validation
To validate our risk prediction framework, we conducted comprehensive simulation tests using the Project Chrono open-source physics engine  (?), which has been systematically validated against real-world experiments  (?). We created a virtual replica of the NASA Ames LHS-1 testbed by using the ground-truth regolith strength measurements collected by the Traveler leg penetrometer to parameterize the simulation environment. Specifically, we computed the average penetration resistance value for each 1 m $\times$ 1 m cell and mapped these values to corresponding elastic and damping parameters for the SCM (Soil Contact Model) terrain representation used in the simulation  (?) (see Appendix Appendix A: Simulation Parameter Mapping for the complete stiffness parameter matrix and mapping methodology). This spatial parameterization creates an 18 m $\times$ 12 m grid where each cell corresponds to a 3 m $\times$ 3 m region of the physical testbed, with penetration resistance per unit area ranging from 0.28 to 8.38 N/cm³ and reflecting the three distinct terrain types (tamped, raked, and sifted), represented by different colors in the simulation visualization (Fig. 4, bottom section). Using this simulation environment, we tested three critical scenarios that demonstrate the importance of payload-aware risk assessment.
@@ 2.3.1 Site selection and field workflow
To demonstrate the advantages of using a scout robot to support scientific target selection and safe rover traversal, we deployed our scout–rover cooperation system at White Sands National Park. White Sands is an active gypsum dune field within a playa–lake basin that displays a rich variety of sedimentary features and surface textures, including loose sand, partially lithified sandstones, and salt and microbial crusts. Features at this site closely resemble the terrains encountered at Meridiani Planum by the NASA Opportunity rover (?, ?, ?), and the existing robust body of work characterizing this dune field facilitates our investigations  (?, ?, ?).
@@ 2.3.3 Terrain-aware rover navigation based on scouting measurements
The integration of risk maps and selected scientific interest goals provides rich information for mission planning that allows rovers to safely navigate to scientific targets while avoiding high-risk areas. To demonstrate this capability, we used a reactive potential field planner  (?, ?), in which goal locations corresponding to high scientific reward generate attractive forces that pull rovers closer and the boundaries of high-risk regions generate repulsive forces that push rovers away from hazardous terrain. The path that the rover takes is then dictated by the sum of attractive and repulsive forces acting on the rover at each point in time. This approach ensures that the rover navigates along a path that achieves scientific discovery goals while minimizing the probability of immobilization. Note that for the purposes of path planning, the rover was modeled as a point mass with a bounded acceleration.
@@ 4.4 Regolith strength map reconstruction
To predict and map the terrain properties over a continuous 2D area, we use Gaussian Process Regression  (?) (GPR), also known as kriging, thereby modeling the underlying terrain property distribution (Fig. 7G, H). By updating the predicted distributions for terrain property map and fit error heuristic map with each step, we provide real-time prediction and quantifiable uncertainty for the terrain properties and the “trustworthiness” across the landscape for planning traversal and science objectives. To use GPR, we define the kernel as
@@ 4.7 Autonomous mission execution through potential-field-based navigation
Figure 2: Experimental setup for terrain-aware legged robot scouting. (A) Experiment site, the SSERVI Regolith Lab testbed located at the NASA Ames Research Center. (B–D) Three terrain-preparation protocols with strength from high to low: tamped (high strength), raked (medium strength), and sifted (low strength). (E)(F) The desired terrain strength map, prepared using the three protocols. (G) Ground truth terrain strength measurements are collected using a robotic penetrometer across the testbed. (H) A legged scout robot performs terrain scouting using proprioception from every step along its path. (I) LiDAR scan after terrain preparation. (J) Penetration resistance measurements from the robotic penetrometer across the testbed. (K) The penetration resistance distribution from terrain grid cells prepared using the three protocols.
@@
Figure 5: Workflow for safe rover mission planning integrating satellite imagery, in-situ measurements, and science-interest-driven goal selection. (A) Overview of the mission plan showing a top-down view of the operational area and the tentative traversal region (purple rectangle). (B) LiDAR-based reconstruction of the traversal region with elevation data (1184–1189 m). (C) LiDAR-based inclination measurements, and initial identification of candidate rover routes. (D) Field interface enabling real-time visualization of terrain strength and science goal selection: (a) estimated toe force over time; (b) toe force during the penetration phase; (c) model fitness; (d) penetration resistance measured from each step; and (e) interpolated penetration resistance map generated via Gaussian Process regression. (E) Field snapshot after Spirit robot scouting. (F) Final penetration resistance map of the scouted region.